OpenStudy (anonymous):

x= 9cos(theta) - 2cos^2 (theta) y= 9sin(theta) - 2sin(theta)cos(theta) find the slope for some reason i wanted to substitute something in for cos^2(theta) like 1/2 cos(2theta) + 1/2. for some reason, when i do this substitution i get a different slope. please help

OpenStudy (anonymous):

\[x = 9\cos(\theta) - 2 \cos^2(\theta)\] \[y = 9\sin(\theta) - 2\sin(\theta)\cos(\theta)\] Right?

OpenStudy (anonymous):

yes